The area of a circular region is numerically equal to four times the circumference of the circle. Find the length of a radius of the circle.
Answer by stanbon(75887)

The area of a circular region is numerically equal to four times the circumference of the circle. Find the length of a radius of the circle.
---------------
Equation:
pi*r^2 = 4(2pi*r)
-----
r^2 = 8r
-----
r(r-8) = 0
Positive solution:
radius = 8
===================
Cheers,
Stan H.
===================